How many elementary operations are used in the following algorithm?
Step 1 Set S = a, k = 1, and t = a.
Step 2 while k < n
(a) Replace t with t + d
(b) Replace S with S + t
(c) Replace k with k + 1
Endwhile
Step 3 Print S
1
Expert's answer
2011-04-04T11:31:22-0400
Step 1: 3 operations (3 "=");
Step 2: 6(n-1) + n = 7(n-1) + 1& operations ((n-1)*(3 "=", 3 "+") and& n checkings if k<n).
Step 3 : print() is not elementary operation.

Thus:& 3 + 7(n-1)+1& = 7(n-1) + 4.